OpenStudy (anonymous):

solve this polynomial equation: 2x^4 - 10x^2 = 72

OpenStudy (solomonzelman):

let \(\large\color{black}{ x^2 =a }\)

OpenStudy (solomonzelman):

re-write your equation in terms of a, and solve it.

OpenStudy (triciaal):

you can also factor out 2 to work with smaller numbers

OpenStudy (anonymous):

that's not how my math teacher taught us how to do it so i dont know what you mean. for this equation, we'd have to set it to 0 so we'd have to subtract 72 from both sides

OpenStudy (solomonzelman):

subtracting 72 from both sides, \(\large\color{black}{ 2x^4 - 10x^2-72\color{red}{-72} = 0\color{red}{-72} }\) this will get you, \(\large\color{black}{ 2x^4 - 10x^2-72 = 0 }\) then say, let \(\large\color{magenta}{ x^2=a }\) lets do our substitution, \(\large\color{black}{ 2x^4 - 10x^2-72 = 0 }\) \(\large\color{black}{ 2(x^2)^2 - 10x^2-72 = 0 }\) \(\large\color{black}{ 2(\color{magenta}{x^2})^2 - 10\color{magenta}{x^2}-72 = 0 }\) \(\large\color{black}{ 2(\color{magenta}{a})^2 - 10\color{magenta}{a}-72 = 0 }\)

OpenStudy (solomonzelman):

So, you get, \(\large\color{black}{ 2a^2 - 10a-72 = 0 }\)

OpenStudy (solomonzelman):

Can you solve this equation?

OpenStudy (anonymous):

honestly, I don't know what to do. I see it and my mind is blank. my dad is currently trying to help me and he's usually really good at math. he doesnt know what the heck to do

OpenStudy (solomonzelman):

You, right now, have to worry just about solving this, \(\Large\color{black}{2a^2-10a-72=0 }\)

OpenStudy (solomonzelman):

\(\Large\color{black}{2(a^2-5a-36)=0 }\) \(\Large\color{black}{a^2-5a-36=0 }\) So far so good? then, I'll give you a hint, \(\Large\color{black}{36=\color{red}{9}\times \color{darkgoldenrod}{4} }\) and \(\Large\color{black}{\color{red}{9}-\color{darkgoldenrod}{4}=5 }\) use this to think how to factor your equation.

OpenStudy (anonymous):

this is what I did for the equation, which was on our test. we're allowed to do test corrections and I got 4 points off, and I don't know why. we also had to find the x values and I got -4 and 9 for it. what did I do wrong?

OpenStudy (solomonzelman):

ohh, these are not the `x` values, you did, \(\large\color{black}{ a^2-5a-36=0 }\) \(\large\color{black}{(a-9)(a+4)=0 }\) \(\large\color{black}{a=-4,~~9 }\) right?

OpenStudy (solomonzelman):

but, we initially said, that \(\large\color{magenta}{a=x^2 }\) , so there was an additional step. \(\large\color{black}{x^2=9 }\) or \(\large\color{black}{x^2=-4 }\)

OpenStudy (anonymous):

yes that's what I did. so would the x's be 3 and -2?

OpenStudy (solomonzelman):

No.

OpenStudy (solomonzelman):

Have you learned about imaginary number, \(\large\color{black}{i }\) ?

OpenStudy (anonymous):

yes i have

OpenStudy (solomonzelman):

Okay, I can see you haven't... \(\large\color{black}{x^2=-4 }\) will have 2 imaginary solutions, \(\large\color{black}{x=\pm2i }\), but no real solutions.

OpenStudy (solomonzelman):

ohh, you have... sorry.

OpenStudy (solomonzelman):

So you can see why I am getting these soltuions? right?

OpenStudy (anonymous):

yeah I see. and 9 would then become x= \[\pm 3i\]

OpenStudy (solomonzelman):

NO.

OpenStudy (anonymous):

calm down, please. I am trying to understand. not everybody is a math wiz

OpenStudy (solomonzelman):

you have, \(\large\color{black}{x^2=9 }\) , and not, \(\large\color{black}{x^2=\color{red}{-}9 }\)

OpenStudy (solomonzelman):

I just said, no. I wasn't harsh, or anything, was I?

OpenStudy (anonymous):

it kinda seemed like it. but don't worry about it. what would we do with 9 then?

OpenStudy (anonymous):

oh I forgot to say this before. I got -4 and 9 because those are what you'd need to get 4 and -9 to equal 0. like for (x-2)(x+2)(x+3), x would equal \[\pm2\] and -3.

OpenStudy (triciaal):

@mcaMohawk @SolomonZelman was not being mean or harsh etc. . He just needs you to stay focused so you can really get it. you can be a math whiz! In your explanation you mean you got -4 and 9 because that's what you need to get -5